select on card from a standard deck. What is the probability of selecting a four, seven or a heart?
Found 2 solutions by stanbon, gonzo:
Answer by stanbon(75887)

select on card from a standard deck. What is the probability of selecting a four, seven or a heart?
-------------------------
# of fours = 4 including a heart
# of sevens = 4 including a heart
# of hearts = 13 including a seven and a heart
-------------------------------
P(4, or 7 or heart) = (3+3+13)/52 = 19/52
=====================
Cheers,
Stan H.

Answer by gonzo(654)

there are 52 cards in a normal set (not including the joker)
4 suits * 13 cards each suit = 52
each suit contains:
number 2 through 10 + j + q + k + a
-----
there are 4 number 4's in a deck.
there are 4 number 7's in a deck.
there are 13 hearts in a deck.
-----
one of the 4's is a heart.
one of the 7's is a heart.
-----
there are 11 hearts in a deck that are not also a 4 and are not also a 7.
-----
total number of cards you can pick that will get you what you want are
4 + 4 + 11
-----
probability of getting what you want is 11/52
-----